# Tensor Product Attention Is All You Need

### v6

@@ 3.3 KV Caching and Memory Reduction
Table 1: Comparison of different attention mechanisms. Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively. Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in the Appendix G. For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
@@
MHA $2hd_{h}$ $4d_{\text{model}}^{2}$ $h$ $h$
MQA $2d_{h}$ $(2+2/h)d_{\text{model}}^{2}$ $h$ 11
GQA $2gd_{h}$ $(2+2g/h)d_{\text{model}}^{2}$ $h$ $g$
MLA $d_{c}+d_{h}^{R}$ $d_{c}^{\prime}(d_{\text{model}}+hd_{h}+hd_{h}^{R})$ $d_{\text{model}}d_{h}^{R}+d_{c}(d_{\text{model}}+2hd_{h})$ $h$ $h$
@@ 4.3 Model Architectures
We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention). Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilizes the overall architecture of LLaMA (touvron2023llama) while changing the self-attention block to our TPA-based version. The feed-forward network (FFN) adopts a SwiGLU layer, as in (shazeer2020glu; touvron2023llama).
Rotary Positional Embedding (RoPE). As discussed in Section 3.2, RoPE (su2024roformer) is applied to the $\mathbf{Q}$ and $\mathbf{K}$. Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see Equation (3.5) and Theorem 3.1.
SwiGLU Feed-Forward Network. Following shazeer2020glu; touvron2023llama, our T6 uses a SwiGLU-based Feed-Forward Network (FFN): $\operatorname{FFN}(\mathbf{x})=\bigl[\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr]\,\bm{W}_{3}$, where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters. Note that other activation functions can also be used.
@@ 5 FlashTPA Decoding Algorithm
Figure 2: Data flow diagram for FlashTPA Decoding. Rectangles represent tensors (blue for inputs, yellow for intermediates, red for final output), circles with $\sum$ or $\odot$ denote Einstein summation contractions or element-wise products respectively, and the green rounded rectangle is the softmax operation. Shapes are shown for a single query ($N=1$) interacting with $M$ cached items. $H$ is the number of heads, $R_{Q}$ is the query rank, and $D,E$ are respective feature dimensions for the $\mathbf{B}_{Q}/\bm{b}^{K}_{\text{cache}}$ and $\bm{b}^{V}_{\text{cache}}$ factors. Scaling factors in softmax are omitted for visual clarity.
@@ 6.1 Language Modeling Tasks
All experiments reported in this paper are implemented based on the nanoGPT codebase (Karpathy2022), and we pretrain our models using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu). The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation. We compare T6 against the baseline Llama architecture (touvron2023llama) with SwiGLU activation (shazeer2020glu) and RoPE embeddings (su2024roformer), as well as Llama variants that replace Multi-Head Attention (MHA; vaswani2017attention) with Multi-Query Attention (MQA; shazeer2019fast), Grouped Query Attention (GQA; ainslie2023gqa), or Multi-head Latent Attention (MLA; liu2024deepseek). In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer. We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B). We pretrain all models with 50B tokens using 1 epoch. Details on architecture hyperparameters and training hardware are shown in Appendix H.1.
Training & Validation Curves. Figure 4 compares validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B. Training loss curves are provided in Appendix Figure 3. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix G) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final validation losses. For instance, in Figure 4(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (liu2024deepseek) (blue curves) generally trains more slowly and yields higher validation losses.
@@
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (ARC), BoolQ (BoolQ), HellaSwag (hellaswag), OBQA (OpenBookQA), PIQA (PIQA), WinoGrande (WinoGrande), and MMLU (MMLU), using the lm-evaluation-harness codebase (eval-harness). For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy. Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3. Zero-shot evaluation of small and XL models are displayed in Tables 11 and 12 in the appendix. Moreover, we also present 2-shot evaluation results in Tables 131415 and 16 in the appendix.
@@ 6.2 Experimental Results on FlashTPA Decoding
This section presents an evaluation of FlashTPA’s decoding time in comparison to several other optimized attention mechanisms. We benchmark FlashTPA against FlashMHA (shah2024flashattention), FlashGQA, FlashMQA, and FlashMLA (flashmla2025). It is important to note that our current FlashTPA implementation utilizes Triton (tillet2019triton). While the compared methods are typically available as highly optimized CUDA kernels, these experiments provide initial insights into FlashTPA’s potential. Development of a CUDA-based FlashTPA kernel is ongoing and is expected to yield further performance improvements.
@@ Acknowledgements
We thank the anonymous reviewers and area chair for their helpful comments. We acknowledge the compute credits provided by Fetch.ai.
[…]
@@ Appendix A Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
$\mathbf{A}_{Q}\in\mathbb{R}^{B\times T_{q}\times h\times R_{Q}},\quad\mathbf{B}_{Q}\in\mathbb{R}^{B\times T_{q}\times R_{Q}\times D},\qquad\mathbf{A}_{K}\in\mathbb{R}^{B\times T_{k}\times h\times R_{K}},\quad\mathbf{B}_{K}\in\mathbb{R}^{B\times T_{k}\times R_{K}\times D},$
$\mathbf{A}_{V}\in\mathbb{R}^{B\times T_{k}\times h\times R_{V}},\qquad\mathbf{B}_{V}\in\mathbb{R}^{B\times T_{k}\times R_{V}\times E}.$
@@ A.1 Direct computation in factor space
$\bigl[\mathbf{Q}^{(h)}(\mathbf{K}^{(h)})^{\top}\bigr]_{q,k}\;=\;\frac{1}{R_{Q}R_{K}}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\,a^{(r)}_{q,h}(\mathbf{x}_{q})\,a^{(s)}_{k,h}(\mathbf{x}_{k})\;\big\langle\mathbf{b}_{q}^{(r)}(\mathbf{x}_{q}),\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{k})\big\rangle,$ (A.1)
and for values (with $s_{V}{=}1/R_{V}$), $\mathbf{V}^{(h)}_{k}=\frac{1}{R_{V}}\sum_{u=1}^{R_{V}}a^{(u)}_{k,h}(\mathbf{x}_{k})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{k})$. The per-head attention output at query position $q$ is then $\sum\nolimits_{k}\operatorname{softmax}\!\big(\frac{1}{\sqrt{D}}[\mathbf{Q}^{(h)}(\mathbf{K}^{(h)})^{\top}]_{q,:}\big)_{k}\,\mathbf{V}^{(h)}_{k}$.
@@
$\mathcal{L}[b,h,q,k]$ $=\mathcal{S}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\mathbf{A}_{Q}[b,q,h,r]\;\mathbf{A}_{K}[b,k,h,s]\;P[b,q,k,r,s],$ (A.2)
$\alpha[b,h,q,k]$ $=\operatorname{Softmax}_{k}\!\left(\mathcal{L}[b,h,q,k]+\log\mathsf{mask}[q,k]\right),$
$\mathbf{O}[b,h,q,e]$ $=s_{V}\sum_{k=1}^{T_{k}}\sum_{u=1}^{R_{V}}\alpha[b,h,q,k]\;\mathbf{A}_{V}[b,k,h,u]\;\mathbf{B}_{V}[b,k,u,e].$ (A.3)
Here $\mathsf{mask}[q,k]\!\in\!\{0,-\infty\}$ enforces causality and padding. Eqs. (A.2)–(A.3) make explicit that (i) feature-space dot products $P$ are head-shared, and (ii) all $1/R_{(\cdot)}$ scalings can be folded into a single multiplicative factor per phase without affecting numerical stability (softmax shift invariance).
@@ A.4 Inference-time decoding cost across mechanisms
  • Attention: The query vectors interact with a K/V cache of size $T$, costing $\Theta(MHd_{h})$.
  • Total MHA: The complexity is $\Theta(d_{\text{model}}Hd_{h}+MHd_{h})$.
@@
  • Total MQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2)+MHd_{h})$.
Grouped-Query Attention (GQA) uses $H$ query heads and $N_{g}$ Key/Value head groups ($H_{kv}=N_{g}$). The arithmetic complexity is also identical to MHA.
  • Projection: Projecting for $H$ query heads and $N_{g}$ K/V head groups costs $\Theta(d_{\text{model}}(Hd_{h}+2N_{g}d_{h}))$.
@@
  • Total GQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2N_{g})+MHd_{h})$.
@@
Multi-Head Latent Attention (MLA), as described in Appendix F.3, uses $H$ heads. For each head, the key $\mathbf{K}_{i}$ has dimension $d_{k}^{\prime}=d_{c}+d_{h}^{R}$, and the value $\mathbf{V}_{i}^{C}$ has dimension $d_{v}^{\prime}=d_{c}$.
  • Projection: Projecting $\mathbf{x}_{M}$ to get compressed and RoPE components for Q, K, and V costs $\Theta(d_{\text{model}}((d_{c}+d_{h}^{R})\cdot h+d_{c}+d_{h}^{R}))$.
  • Attention: The query vectors interact with a K/V cache of size $M$, costing $\Theta(MH(2d_{c}+d_{h}^{R}))$.
  • Total MLA: The complexity is $\Theta(d_{\text{model}}(d^{\prime}_{c}+2d_{c}+2d_{h}^{R})+MH(2d_{c}+d_{h}^{R}))$.
TPA. We use the FlashTPA Decoding algorithm (Algorithm 2) for FLOPs analysis, with $N=1$ query token, $T$ cached items, $D$ as feature dimension for $\mathbf{B}_{Q}/\bm{b}^{K}$ (typically $d_{h}$), and $E$ for $\bm{b}^{V}$ (typically $d_{h}$). For ranks $(R_{Q},R_{K},R_{V})$:
  • Projection: Projecting $\mathbf{x}_{M}$ to all Q, K, V factors costs $\Theta(d_{\text{model}}((R_{Q}+R_{K}+R_{V})(H+d_{h})))$.
  • Attention: The decoding algorithm’s interaction with the cache costs $\Theta(M[R_{K}(R_{Q}D+HR_{Q}+H)+R_{V}H(1+E)])$.
  • Total for TPA decoding: $\Theta(d_{\text{model}}(R_{Q}+R_{K}+R_{V})(H+d_{h})+M[R_{K}(R_{Q}D+HR_{Q}+H)+R_{V}H(1+E)])$.
@@
GQA ($N_{g}=4$, 9.4M parameters):
$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=2048\cdot 64\cdot(2\cdot 32+2\cdot 4)\approx 9.4\times 10^{6}$
$\text{Projection}=d_{\text{model}}(H+2N_{g})d_{h}=2048\cdot(32+8)\cdot 64\approx 5.2\times 10^{6}$
@@
$\text{Projection}=d_{\text{model}}((d_{c}+d_{h}^{R})\cdot h+d_{c}+d_{h}^{R})=2048\cdot((256+32)\cdot 32+256+32)\approx 19.5\times 10^{6}$
@@
$\text{Attention}=M\cdot[1(1568)+1(2080)]=3648M$
@@
$\text{Attention}=M\cdot[2(1568)+2(2080)]=7296M$
@@
$\text{Attention}=M\cdot[1(800)+1(2080)]=2880M$
@@
$\text{Attention}=M\cdot[2(800)+2(2080)]=5760M$
@@
GQA ($N_{g}=4$, 37.7M parameters):
$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=4096\cdot 128\cdot(2\cdot 32+2\cdot 4)\approx 37.7\times 10^{6}$
@@
$\text{Projection}=4096\cdot((512+64)\cdot 32+512+64)\approx 77.9\times 10^{6}$
@@
$\text{Attention}=M\cdot[1(2592)+1(4128)]=6720M$
@@
$\text{Attention}=M\cdot[2(2592)+2(4128)]=13440M$
@@
$\text{Attention}=M\cdot[1(1312)+1(4128)]=5440M$
@@
$\text{Attention}=M\cdot[2(1312)+2(4128)]=10880M$

### v7

@@ 3.3 KV Caching and Memory Reduction
Table 1: Comparison of different attention mechanisms. Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively. Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in the Appendix G. For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively. The MLA parameter count includes the output projection.
@@
MHA $2hd_{h}$ $4d_{\text{model}}\,h\,d_{h}$ $h$ $h$
MQA $2d_{h}$ $2d_{\text{model}}\,d_{h}\,(h+1)$ $h$ 11
GQA $2Gd_{h}$ $2d_{\text{model}}\,d_{h}\,(h+G)$ $h$ $G$
MLA $d_{c}+d_{h}^{R}$ $d_{c}^{\prime}(d_{\text{model}}+hd_{h}+hd_{h}^{R})$ $d_{c}(d_{\text{model}}+2hd_{h}) +$$d_{\text{model}}(hd_{h}+d_{h}^{R})$ $h$ $h$
@@ 4.3 Model Architectures
We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention). Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilizes the overall architecture of LLaMA (Touvron et al., 2023) while changing the self-attention block to our TPA-based version. The feed-forward network (FFN) adopts a SwiGLU layer, as in (Shazeer, 2020; Touvron et al., 2023).
Rotary Positional Embedding (RoPE). As discussed in Section 3.2, RoPE (Su et al., 2024) is applied to the $\mathbf{Q}$ and $\mathbf{K}$. Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see Equation (3.5) and Theorem 3.1.
SwiGLU Feed-Forward Network. Following Shazeer (2020); Touvron et al. (2023), our T6 uses a SwiGLU-based Feed-Forward Network (FFN): $\operatorname{FFN}(\mathbf{x})=\bigl[\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr]\,\bm{W}_{3}$, where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters. Note that other activation functions can also be used.
@@ 5 FlashTPA Decoding Algorithm
Figure 2: Data flow diagram for FlashTPA Decoding. Rectangles represent tensors (blue for inputs, yellow for intermediates, red for final output), circles with $\sum$ or $\odot$ denote Einstein summation contractions or element-wise products respectively, and the green rounded rectangle is the softmax operation. Shapes are shown for a single query ($N=1$) interacting with $M$ cached items in the common rank-1 setting $R_{K}=R_{V}=1$. We use a head-first layout $(H,M)$ for logits and attention weights; the cached head factors $\mathbf{a}^{K}_{\text{cache}}$ and $\mathbf{a}^{V}_{\text{cache}}$ are shown transposed relative to their natural token-major layout for readability. $H$ is the number of heads, $R_{Q}$ is the query rank, and $D,E$ are respective feature dimensions for the $\mathbf{B}_{Q}/\bm{b}^{K}_{\text{cache}}$ and $\bm{b}^{V}_{\text{cache}}$ factors. Scaling factors are omitted for visual clarity.
@@ 6.1 Language Modeling Tasks
All experiments reported in this paper are implemented based on the nanoGPT codebase (Karpathy, 2022), and we pretrain our models using the FineWeb-Edu 100B dataset (Lozhkov et al., 2024). The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation. We compare T6 against the baseline Llama architecture (Touvron et al., 2023) with SwiGLU activation (Shazeer, 2020) and RoPE embeddings (Su et al., 2024), as well as Llama variants that replace Multi-Head Attention (MHA; Vaswani et al. (2017)) with Multi-Query Attention (MQA; Shazeer (2019)), Grouped Query Attention (GQA; Ainslie et al. (2023)), or Multi-head Latent Attention (MLA; Liu et al. (2024a)). In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer. We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B). We pretrain all models for 50B tokens (roughly half an epoch over FineWeb-Edu-100B). Details on architecture hyperparameters and training hardware are shown in Appendix H.1.
Training & Validation Curves. Figure 4 compares validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B. Training loss curves are provided in Appendix Figure 3. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix G) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final validation losses. For instance, in Figure 4(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (Liu et al., 2024a) (blue curves) generally trains more slowly and yields higher validation losses.
@@
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020), and MMLU (Hendrycks et al., 2021), using the lm-evaluation-harness codebase (Gao et al., 2024). For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy. Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3. Zero-shot evaluation of small and XL models are displayed in Tables 11 and 12 in the appendix. Moreover, we also present 2-shot evaluation results in Tables 131415 and 16 in the appendix.
@@ 6.2 Experimental Results on FlashTPA Decoding
This section presents an evaluation of FlashTPA’s decoding time in comparison to several other optimized attention mechanisms. We benchmark FlashTPA against FlashMHA (Shah et al., 2024), FlashGQA, FlashMQA, and FlashMLA (Jiashi Li, 2025). It is important to note that our current FlashTPA implementation utilizes Triton (Tillet et al., 2019). While the compared methods are typically available as highly optimized CUDA kernels, these experiments provide initial insights into FlashTPA’s potential. Development of a CUDA-based FlashTPA kernel is ongoing and is expected to yield further performance improvements.
@@ Acknowledgements
We thank the anonymous reviewers and area chairs for their helpful comments. We acknowledge the compute credits provided by Fetch.ai.
@@ Appendix A Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
$\mathbf{A}_{Q}\in\mathbb{R}^{B\times T_{q}\times R_{Q}\times H},\quad\mathbf{B}_{Q}\in\mathbb{R}^{B\times T_{q}\times R_{Q}\times D},\qquad\mathbf{A}_{K}\in\mathbb{R}^{B\times T_{k}\times R_{K}\times H},\quad\mathbf{B}_{K}\in\mathbb{R}^{B\times T_{k}\times R_{K}\times D},$
$\mathbf{A}_{V}\in\mathbb{R}^{B\times T_{k}\times R_{V}\times H},\qquad\mathbf{B}_{V}\in\mathbb{R}^{B\times T_{k}\times R_{V}\times E}.$
@@
Convention. For a single token, the main text defines $\mathbf{A}_{*}(\mathbf{x}_{t})\in\mathbb{R}^{R_{*}\times H}$ and $\mathbf{B}_{*}(\mathbf{x}_{t})\in\mathbb{R}^{R_{*}\times D}$, with $\mathbf{Q}_{t}=\tfrac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$. Accordingly, throughout this appendix we index $\mathbf{A}_{Q}$ as $\mathbf{A}_{Q}[b,q,r,h]$ (rank-major). Some implementations may store $\mathbf{A}_{*}$ transposed as $(H\times R_{*})$ for memory layout, this is equivalent, since all uses contract over the rank index.
@@ A.1 Direct computation in factor space
$\bigl[\mathbf{Q}^{(h)}(\mathbf{K}^{(h)})^{\top}\bigr]_{q,k}\;=\;\frac{1}{R_{Q}R_{K}}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\,a^{Q,(r)}_{q,h}(\mathbf{x}_{q})\,a^{K,(s)}_{k,h}(\mathbf{x}_{k})\;\big\langle\mathbf{b}^{Q,(r)}(\mathbf{x}_{q}),\,\mathbf{b}^{K,(s)}(\mathbf{x}_{k})\big\rangle,$ (A.1)
and for values (with $s_{V}{=}1/R_{V}$), $\mathbf{V}^{(h)}_{k}=\frac{1}{R_{V}}\sum_{u=1}^{R_{V}}a^{V,(u)}_{k,h}(\mathbf{x}_{k})\,\mathbf{b}^{V,(u)}(\mathbf{x}_{k})$. The per-head attention output at query position $q$ is then $\sum\nolimits_{k}\operatorname{softmax}\!\big(\frac{1}{\sqrt{D}}[\mathbf{Q}^{(h)}(\mathbf{K}^{(h)})^{\top}]_{q,:}\big)_{k}\,\mathbf{V}^{(h)}_{k}$.
@@
$\mathcal{L}[b,h,q,k]$ $=\mathcal{S}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\mathbf{A}_{Q}[b,q,r,h]\;\mathbf{A}_{K}[b,k,s,h]\;P[b,q,k,r,s],$ (A.2)
$\alpha[b,h,q,k]$ $=\operatorname{Softmax}_{k}\!\left(\mathcal{L}[b,h,q,k]+\mathsf{mask}[b,q,k]\right),$
$\mathbf{O}[b,h,q,e]$ $=s_{V}\sum_{k=1}^{T_{k}}\sum_{u=1}^{R_{V}}\alpha[b,h,q,k]\;\mathbf{A}_{V}[b,k,u,h]\;\mathbf{B}_{V}[b,k,u,e].$ (A.3)
Here $\mathsf{mask}[b,q,k]\!\in\!\{0,-\infty\}$ is an additive mask in logit space that enforces causality and padding. Eqs. (A.2)–(A.3) make explicit that (i) feature-space dot products $P$ are head-shared, and (ii) the rank normalizations $s_{Q},s_{K},s_{V}$ can be absorbed into the corresponding factor tensors (or into the scalar prefactors) without changing the computed attention output.
@@ A.4 Inference-time decoding cost across mechanisms
  • Attention: Dot products and value aggregation over a cache of length $M$ cost $\Theta(2MHd_{h})$ (ignoring softmax constants).
  • Total MHA: The complexity is $\Theta(d_{\text{model}}Hd_{h}+2MHd_{h})$.
@@
  • Total MQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2)+2MHd_{h})$.
Grouped-Query Attention (GQA) uses $H$ query heads and $G$ Key/Value head groups ($H_{kv}=G$). The arithmetic complexity is also identical to MHA.
  • Projection: Projecting for $H$ query heads and $G$ K/V head groups costs $\Theta(d_{\text{model}}(Hd_{h}+2Gd_{h}))$.
@@
  • Total GQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2G)+2MHd_{h})$.
@@
Multi-Head Latent Attention (MLA), as described in Appendix F.3, uses $H$ heads. Each head’s (up-projected) query/key vectors have dimension $d_{h}+d_{h}^{R}$. During decoding, however, the score computation against the cache can be decomposed into (i) a dot product in the cached latent space $\mathbb{R}^{d_{c}}$ for the content part and (ii) an additional RoPE dot product in $\mathbb{R}^{d_{h}^{R}}$ for the positional part. Concretely, MLA caches $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ per past token $s$, aggregates values in $\mathbb{R}^{d_{c}}$, and then up-projects once per step.
  • Cached state: MLA caches the compressed KV latent $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ and the shared RoPE key component $\mathbf{k}_{s}^{R}\in\mathbb{R}^{d_{h}^{R}}$ per past token $s$.
  • Projection (current token): Computing the query latents and the new cache entry (up to constant factors) costs
    $\Theta\bigl(d_{\text{model}}d_{c}^{\prime}\;+\;d_{c}^{\prime}H(d_{h}+d_{h}^{R})\;+\;d_{\text{model}}(d_{c}+d_{h}^{R})\bigr),$
    corresponding to forming $\mathbf{c}^{Q}$, $\mathbf{Q}^{C}$, $\mathbf{Q}^{R}$, and computing/storing $\mathbf{c}^{KV}$ and $\mathbf{k}^{R}$ for the current token.
  • Attention (cache interaction): Using the identity $\mathbf{q}_{t,i}^{C\top}\mathbf{k}_{s,i}^{C}=(\bm{W}^{UK}_{i}\mathbf{q}_{t,i}^{C})^{\top}\mathbf{c}_{s}^{KV}$, the score against each cached token can be computed via a dot product in $\mathbb{R}^{d_{c}}$ plus the RoPE dot product in $\mathbb{R}^{d_{h}^{R}}$. The latent value can be aggregated in $\mathbb{R}^{d_{c}}$ and then up-projected once. The dominant cache-dependent cost is
    $\Theta\bigl(MH(2d_{c}+d_{h}^{R})\bigr),$
    up to lower-order per-step terms such as $\Theta(Hd_{c}d_{h})$.
  • Total MLA: $\Theta\bigl(d_{\text{model}}d_{c}^{\prime}+d_{c}^{\prime}H(d_{h}+d_{h}^{R})+d_{\text{model}}(d_{c}+d_{h}^{R})+MH(2d_{c}+d_{h}^{R})\bigr)$.
TPA. We use the FlashTPA Decoding algorithm (Algorithm 2) for FLOPs analysis, with $N=1$ query token, $M$ cached items, $D$ as feature dimension for $\mathbf{B}_{Q}/\bm{b}^{K}$ (typically $d_{h}$), and $E$ for $\bm{b}^{V}$ (typically $d_{h}$). For ranks $(R_{Q},R_{K},R_{V})$:
  • Projection: Projecting the current token $\mathbf{x}_{T}$ to all Q/K/V factors costs $\Theta\bigl(d_{\text{model}}(R_{Q}{+}R_{K}{+}R_{V})(H{+}d_{h})\bigr)$.
  • Attention (cache interaction): Using Algorithm 2 with cache length $M$, the dominant cache-dependent FLOPs are
    $\Theta\bigl(M\,(R_{Q}R_{K}D\;+\;HR_{Q}R_{K}\;+\;HR_{V}E)\bigr),$
    up to lower-order terms (masking/element-wise products and online-softmax bookkeeping).
  • Total for TPA decoding: $\Theta\bigl(d_{\text{model}}(R_{Q}{+}R_{K}{+}R_{V})(H{+}d_{h})\;+\;M(R_{Q}R_{K}D+HR_{Q}R_{K}+HR_{V}E)\bigr)$.
@@
GQA ($G=4$, 9.4M parameters):
$\text{Parameters}=d_{\text{model}}d_{h}(2H+2G)=2048\cdot 64\cdot(2\cdot 32+2\cdot 4)\approx 9.4\times 10^{6}$
$\text{Projection}=d_{\text{model}}(H+2G)d_{h}=2048\cdot(32+8)\cdot 64\approx 5.2\times 10^{6}$
@@
$\text{Projection}\approx d_{\text{model}}d_{c}^{\prime}+d_{c}^{\prime}H(d_{h}+d_{h}^{R})+d_{\text{model}}(d_{c}+d_{h}^{R})+Hd_{c}d_{h}$
$\hskip 42.19432pt=2048\cdot 768+768\cdot 32\cdot(64+32)+2048\cdot(256+32)+32\cdot 256\cdot 64$
$\hskip 42.19432pt\approx 5.0\times 10^{6}$
@@
$\text{Attention}=M\cdot[1(1536)+1(2048)]=3584M$
@@
$\text{Attention}=M\cdot[2(1536)+2(2048)]=7168M$
@@
$\text{Attention}=M\cdot[1(768)+1(2048)]=2816M$
@@
$\text{Attention}=M\cdot[2(768)+2(2048)]=5632M$
@@
GQA ($G=4$, 37.7M parameters):
$\text{Parameters}=d_{\text{model}}d_{h}(2H+2G)=4096\cdot 128\cdot(2\cdot 32+2\cdot 4)\approx 37.7\times 10^{6}$
@@
$\text{Projection}\approx d_{\text{model}}d_{c}^{\prime}+d_{c}^{\prime}H(d_{h}+d_{h}^{R})+d_{\text{model}}(d_{c}+d_{h}^{R})+Hd_{c}d_{h}$
$\hskip 42.19432pt=4096\cdot 1536+1536\cdot 32\cdot(128+64)+4096\cdot(512+64)+32\cdot 512\cdot 128$
$\hskip 42.19432pt\approx 20.2\times 10^{6}$
@@
$\text{Attention}=M\cdot[1(2560)+1(4096)]=6656M$
@@
$\text{Attention}=M\cdot[2(2560)+2(4096)]=13312M$
@@
$\text{Attention}=M\cdot[1(1280)+1(4096)]=5376M$
@@
$\text{Attention}=M\cdot[2(1280)+2(4096)]=10752M$